# Isolated Causal Effects of Natural Language

### v1

@@ 6 RELATED WORK
Causal effects of text.
Our work on isolated causal effects is situated within a recent literature on estimating text-based causal effects. Egami et al., (2022) describe a conceptual codebook framework for causal inference using text. Building on this idea, Fong and Grimmer, (2023) conduct randomized text experiments where texts are programmatically generated from pre-specified attributes. Lin et al., (2023) further propose a method for transporting natural (i.e., non-isolated) causal effects from randomized text experiments to potentially non-randomized target distributions.
Most directly related to our work are several methods for estimating isolated effects of natural language using specific language representation pipelines. Though these works do not explicitly distinguish isolated and natural effects, their estimands are defined such that they are isolated, and so we view these methods as complementary to ours. Pryzant et al., (2021) estimate the effect of a proxy linguistic attribute from observational data, where all other language information is represented by an embedding from a transformer trained to capture confounding. Dhawan et al., (2024) estimate effects from observational data using the LLM prompting approach we describe earlier. Finally, a very recent paper by Imai and Nakamura, (2024) estimates text effects via a randomized experiment in which LLM-generated texts are shown to human respondents; text representations can then be extracted directly from the generating LLM.
OVB and representation quality.
The diversity of language representations that can be used in text-based causal inference highlights the strong need for a way to analyze the quality of representations and effect estimates. Our OVB-based metrics provide a way to do this flexibly across any data setting, which is important for real-world data where the ground truth is not known. Our metrics draw directly on the Chernozhukov et al., (2024) operationalization of OVB, which in turn builds on a history of foundational work on OVB (Goldberger,, 1991; Frank,, 2000; Angrist and Pischke,, 2009; Oster,, 2019; Cinelli and Hazlett,, 2019).
Noting the importance of covariate representation in causal inference, Clivio et al., 2024b have proposed learning representations that minimize information loss when balancing covariates. Subsequent work shows that this information loss can be used jointly with a confounding error to improve overlap for covariate representations (Clivio et al., 2024a, ). While these methods are not developed specifically for text, the parallels between general covariate representation and language representation are evident.
@@ 7 CONCLUSION
In this paper, we propose a framework for estimating the isolated causal effect of a focal language attribute from natural language. Estimating isolated effects is challenging because it requires us to be able to model the non-focal language. We introduce metrics for evaluating the quality of non-focal language approximations and their corresponding isolated effect estimates along the axes of fidelity and overlap. We demonstrate the ability of our framework to correctly recover isolated effects across a variety of language-encoded interventions, and we explore how the way we approximate non-focal language impacts fidelity, overlap, and the effect estimates themselves. Our results point to a compelling line of future work on learning representations—perhaps not only of language—that optimize the fidelity-overlap tradeoff, making them explicitly suitable for the task of causal inference.
Supplementary Materials
1 DERIVATIONS AND PROOFS
1.1 ATE and ATT
1.1.1 Identifying the estimand
@@
$=\mathbb{E}_{Y(\cdot)\sim\mathcal{G}}\left[\mathbb{E}_{X}\left[\sum_{x\in\mathcal{X}}\frac{P^{*}(a^{\mathsf{c}}(x))}{P(a^{\mathsf{c}}(x),a(x)=a^{\prime})}(Y(a^{\prime},a^{\mathsf{c}}(x))-g(a^{\prime},a^{\mathsf{c}}(x)))\mathbbm{1}\{a(x)=a^{\prime}\}\mathbbm{1}\{a^{\mathsf{c}}(X)=a^{\mathsf{c}}(x),a(X)=a(x)\}\right]\right]$
$=\mathbb{E}_{Y(\cdot)\sim\mathcal{G}}\left[\sum_{x\in\mathcal{X}}\frac{P^{*}(a^{\mathsf{c}}(x))}{P(a^{\mathsf{c}}(x),a(x)=a^{\prime})}(Y(a^{\prime},a^{\mathsf{c}}(x))-g(a^{\prime},a^{\mathsf{c}}(x)))\mathbb{E}_{X}\left[\mathbbm{1}\{a(x)=a^{\prime}\}\mathbbm{1}\{a^{\mathsf{c}}(X)=a^{\mathsf{c}}(x),a(X)=a(x)\}\right]\right]$
@@
1.1.2 $\gamma$ for two special cases
@@ 1.1.3 A general $P^{*}$
$=\frac{{\color[rgb]{.5,0,.5}P(a^{\mathsf{c}}(X)|C=T)}}{{\color[rgb]{0,.5,.5}P(a^{\mathsf{c}}(X)|C=S)}{\color[rgb]{0,0,1}P(a(X)=a^{\prime}|a^{\mathsf{c}}(X),C=S)}}$
$=\frac{{\color[rgb]{.5,0,.5}P(C=T|a^{\mathsf{c}}(X))P(a^{\mathsf{c}}(X))}{\color[rgb]{0,.5,.5}P(C=S)}}{{\color[rgb]{.5,0,.5}P(C=T)}{\color[rgb]{0,.5,.5}P(C=S|a^{\mathsf{c}}(X))P(a^{\mathsf{c}}(X))}{\color[rgb]{0,0,1}P(a(X)=a^{\prime}|a^{\mathsf{c}}(X),C=S)}}$
$=\frac{{\color[rgb]{0,.5,.5}P(C=S)}}{{\color[rgb]{.5,0,.5}P(C=T)}}\times\frac{{\color[rgb]{.5,0,.5}P(C=T|a^{\mathsf{c}}(X))}}{{\color[rgb]{0,.5,.5}P(C=S|a^{\mathsf{c}}(X))}}\times\frac{1}{{\color[rgb]{0,0,1}P(a(X)=a^{\prime}|a^{\mathsf{c}}(X),C=S)}}$
@@
1.1.4 Unbiasedness of $\widehat{\tau}_{DR}$ given one correct model
@@
$\mathbb{E}_{Y\sim P_{y}}$ $[\mathbb{E}_{X,X^{*}}[\widehat{\tau}_{DR}]]=\mathbb{E}_{Y\sim P_{y}}\Big{[}E_{X,X^{*}}\Big{[}\frac{1}{m}\sum_{j=1}^{m}\Big{[}\widehat{g}(1,a_{s}^{\mathsf{c}}(X^{*}_{j}))-\widehat{g}(0,a_{s}^{\mathsf{c}}(X^{*}_{j}))\Big{]}+\frac{1}{n}\sum_{i=1}^{n}\widehat{\gamma}(a(X_{i}),a_{s}^{\mathsf{c}}(X_{i}))(Y_{i}-\widehat{g}(a(X_{i}),a_{s}^{\mathsf{c}}(X_{i})))\Big{]}\Big{]}$
$=\frac{1}{m}\sum_{j=1}^{m}\mathbb{E}_{Y\sim P_{y}}[\mathbb{E}_{X^{*}}\left[\widehat{g}(1,a_{s}^{\mathsf{c}}(X^{*}_{j}))-\widehat{g}(0,a_{s}^{\mathsf{c}}(X^{*}_{j}))\right]]+\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}_{Y\sim P_{y}}[\mathbb{E}_{X}[\widehat{\gamma}(a(X_{i}),a_{s}^{\mathsf{c}}(X_{i}))(Y_{i}-\widehat{g}(a(X_{i}),a_{s}^{\mathsf{c}}(X_{i})))]]$
$=\frac{1}{m}\sum_{j=1}^{m}\mathbb{E}_{X^{*}}\left[\widehat{g}(1,a_{s}^{\mathsf{c}}(X^{*}_{j}))-\widehat{g}(0,a_{s}^{\mathsf{c}}(X^{*}_{j}))\right]+\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}_{Y\sim P_{y}}[\mathbb{E}_{X}[\mathbbm{1}\{a(X_{i})=1\}\widehat{\gamma}(1,a_{s}^{\mathsf{c}}(X_{i}))(Y_{i}-\widehat{g}(1,a_{s}^{\mathsf{c}}(X_{i})))$
$\hskip 250.00038pt+\mathbbm{1}\{a(X_{i})=0\}\widehat{\gamma}(0,a_{s}^{\mathsf{c}}(X_{i}))(Y_{i}-\widehat{g}(0,a_{s}^{\mathsf{c}}(X_{i})))]]$
@@
$\quad-\Bigg{(}\frac{1}{m}\sum_{j=1}^{m}\mathbb{E}_{X^{*}}[\widehat{g}(0,a_{s}^{\mathsf{c}}(X^{*}_{j}))]+\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}_{Y\sim P_{y}}\left[\mathbb{E}_{X^{*}}\left[\mathbbm{1}\{a(X_{i})=0\}\frac{P^{*}(a^{\mathsf{c}}(X_{i}))}{\widehat{P}(a_{s}^{\mathsf{c}}(X_{i}),a(X_{i})=0)}(Y_{i}-\widehat{g}(0,a_{s}^{\mathsf{c}}(X_{i}))\right]\right]\Bigg{)}$
@@
$=\mathbb{E}_{Y(\cdot)\sim\mathcal{G}}\left[\mathbb{E}_{X}\left[\sum_{x\in\mathcal{X}}\frac{P^{*}(a^{\mathsf{c}}(x))}{P(a^{\mathsf{c}}(x),a(x)=a^{\prime})}(Y(a^{\prime},a^{\mathsf{c}}(x))-\widehat{g}(a^{\prime},a_{s}^{\mathsf{c}}(x)))\mathbbm{1}\{a(x)=a^{\prime}\}\mathbbm{1}\{a^{\mathsf{c}}(X)=a^{\mathsf{c}}(x),a(X)=a(x)\}\right]\right]$
@@
$=\mathbb{E}_{Y(\cdot)\sim\mathcal{G}}\left[\mathbb{E}_{X}\left[\sum_{x\in\mathcal{X}}\mathbbm{1}\{a(x)=a^{\prime}\}\frac{P^{*}(a^{\mathsf{c}}(x))}{\widehat{P}(a_{s}^{\mathsf{c}}(x),a(x)=a^{\prime})}(Y(a^{\prime},a^{\mathsf{c}}(x))-g(a^{\prime},a^{\mathsf{c}}(x)))\mathbbm{1}\{a^{\mathsf{c}}(X)=a^{\mathsf{c}}(x),a(X)=a(x)\}\right]\right]$
$=\mathbb{E}_{Y(\cdot)\sim\mathcal{G}}\left[\sum_{x\in\mathcal{X}}\frac{P^{*}(a^{\mathsf{c}}(x))}{\widehat{P}(a_{s}^{\mathsf{c}}(x),a(x)=a^{\prime})}(Y(a^{\prime},a^{\mathsf{c}}(x))-g(a^{\prime},a^{\mathsf{c}}(x)))\mathbb{E}_{X}\left[\mathbbm{1}\{a(x)=a^{\prime}\}\mathbbm{1}\{a^{\mathsf{c}}(X)=a^{\mathsf{c}}(x),a(X)=a(x)\}\right]\right]$
@@
1.1.5 Confidence intervals for $\widehat{\tau}_{DR}$
Consider data $(\widetilde{X}_{1},\dots,\widetilde{X}_{k})=(X_{1},\dots,X_{n},X^{*}_{1},\dots,X^{*}_{m})$, $(\widetilde{Y}_{1},\dots,\widetilde{Y}_{k})=(Y_{1},\dots,Y_{n},0,\dots,0)$. Then following standard procedures for doubly robust estimators (Kennedy,, 2023), the estimator for the closed-form variance of $\widehat{\tau}_{DR}$ is derived from the influence function as follows.

### v3

@@ 6 Related Work
Causal effects of text. Our work on isolated causal effects is situated within a recent literature on estimating text-based causal effects. Egami et al. (2022) describe a conceptual codebook framework for causal inference using text. Building on this, Fong & Grimmer (2023) conduct randomized text experiments where texts are programmatically generated from pre-specified attributes. Lin et al. (2023) further propose a method for transporting natural (i.e., non-isolated) causal effects from randomized text experiments to potentially non-randomized target distributions.
Most directly related to our work are several methods for estimating isolated effects of natural language using specific language representation pipelines. Though these works do not explicitly distinguish isolated and natural effects, their estimands are defined such that they are isolated, and so we view these methods as complementary to ours. Pryzant et al. (2021) estimate the effect of a proxy linguistic attribute from observational data, where all other language information is represented by an embedding from a transformer trained to capture confounding. Dhawan et al. (2024) estimate effects from observational data using the LLM prompting approach we describe earlier. Finally, a recent paper by Imai & Nakamura (2024) estimates text effects via a randomized experiment in which LLM-generated texts are shown to human respondents; text representations can then be extracted directly from the generating LLM.
Omitted variable bias. The diversity of language representations that can be used in text-based causal inference highlights the strong need for a way to understand the quality of representations and effect estimates. Our OVB-based metrics provide a way to do this flexibly across any data setting, which is important for real-world data where the ground truth is not known. Our metrics draw directly on the Chernozhukov et al. (2024) operationalization of OVB, which in turn builds on a history of foundational work on OVB (Goldberger, 1991; Frank, 2000; Angrist & Pischke, 2009; Oster, 2019; Cinelli & Hazlett, 2019).
Noting the importance of covariate representation in causal inference, Clivio et al. (2024b) have proposed learning representations that minimize information loss when balancing covariates, which can help to improve overlap (Clivio et al., 2024a). While these methods are not developed specifically for text, the parallels between general covariate representation and language representation are evident.
@@ 7 Conclusion
In this paper, we propose a framework for estimating the isolated causal effect of a focal language-based intervention. Estimating isolated effects is challenging because it requires us to model not only the focal intervention but also the non-focal language of the text. We introduce measures for assessing the sensitivity of isolated effect estimates to omitted variable bias in their non-focal language approximations along the axes of fidelity and overlap. We demonstrate the ability of our framework to correctly recover isolated effects across multiple language-encoded interventions, and we explore how the way we approximate non-focal language impacts fidelity, overlap, and the robustness of the effect estimates themselves.
Our results point to several avenues for future research. This paper studies a setting in which confounding is contained fully in the text. Though NLP datasets are not often released with external confounding data like information about annotators, it may still be interesting to study the case where external confounding is present and measured. Additionally, in this paper we treat the focal language-encoding function $a(\cdot)$ as an accurate parameterization of the intervention of interest, but if $a(\cdot)$ does need to be estimated, then estimation error can lead to additional bias. Characterizing and counteracting this is important future work. Finally, our findings on OVB and robustness suggest a compelling line of research on learning representations—perhaps not only of language—that optimize the fidelity-overlap tradeoff to minimize omitted variable bias, making them explicitly suitable for the task of causal inference.
Impact Statement
Broader impact. Recent advances in NLP have dramatically increased the availability of language data and models for common users. The resulting proliferation of texts and models has raised potential ethical concerns around factual inaccuracies (Monteith et al., 2024; Zhou et al., 2023), bias (Wan et al., 2023; Ferrara, 2024), and the black-box internals of models (Guidotti et al., 2018; McDermid et al., 2021). These concerns emphasize the growing need to understand the impacts of texts and language models on the readers that consume them. Our work on isolated causal effects builds toward this goal.
Ethical considerations. The empirical analysis contained in this work relies partially on representations from pre-trained large language models, which may encode biases from their training data. Interpretations of causal effects that rely on such representations should consider these biases. We additionally acknowledge the environmental impact of training the language models used in this work.
Acknowledgements
This material is based upon work partially supported by the National Institutes of Health (awards R01MH125740, R01MH132225, R21MH130767, and U01MH136535). Victoria Lin is supported by a Meta Research PhD Fellowship. Any opinions, findings, conclusions, or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the sponsors, and no official endorsement should be inferred.
Appendix A Derivations and Proofs
A.1 ATE and ATT
A.1.1 Identifying the Estimand
@@
$=\mathbb{E}_{Y(\cdot)\sim\mathcal{G}}\Bigg{[}\mathbb{E}_{X}\Bigg{[}\sum_{x\in\mathcal{X}}\frac{P^{*}(a^{\mathsf{c}}(x))}{P(a^{\mathsf{c}}(x),a(x)=a^{\prime})}(Y(a^{\prime},a^{\mathsf{c}}(x))$
$\quad-g(a^{\prime},a^{\mathsf{c}}(x)))\mathbbm{1}\{a(x)=a^{\prime}\}\mathbbm{1}\{a^{\mathsf{c}}(X)=a^{\mathsf{c}}(x),a(X)=a(x)\}\Bigg{]}\Bigg{]}$
$=\mathbb{E}_{Y(\cdot)\sim\mathcal{G}}\Bigg{[}\sum_{x\in\mathcal{X}}\frac{P^{*}(a^{\mathsf{c}}(x))}{P(a^{\mathsf{c}}(x),a(x)=a^{\prime})}(Y(a^{\prime},a^{\mathsf{c}}(x))$
$\quad-g(a^{\prime},a^{\mathsf{c}}(x)))\mathbb{E}_{X}\Bigg{[}\mathbbm{1}\{a(x)=a^{\prime}\}\mathbbm{1}\{a^{\mathsf{c}}(X)=a^{\mathsf{c}}(x),a(X)=a(x)\}\Bigg{]}\Bigg{]}$
@@
A.1.2 $\gamma$ for Two Special Cases
@@ A.1.3 A General $P^{*}$
$=\frac{{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}P(a^{\mathsf{c}}(X)|C=T)}}{{\color[rgb]{0,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{0,.5,.5}P(a^{\mathsf{c}}(X)|C=S)}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}P(a(X)=a^{\prime}|a^{\mathsf{c}}(X),C=S)}}$
$=\frac{{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}P(C=T|a^{\mathsf{c}}(X))P(a^{\mathsf{c}}(X))}{\color[rgb]{0,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{0,.5,.5}P(C=S)}}{{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}P(C=T)}{\color[rgb]{0,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{0,.5,.5}P(C=S|a^{\mathsf{c}}(X))P(a^{\mathsf{c}}(X))}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}P(a(X)=a^{\prime}|a^{\mathsf{c}}(X),C=S)}}$
$=\frac{{\color[rgb]{0,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{0,.5,.5}P(C=S)}}{{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}P(C=T)}}\times\frac{{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}P(C=T|a^{\mathsf{c}}(X))}}{{\color[rgb]{0,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{0,.5,.5}P(C=S|a^{\mathsf{c}}(X))}}\times\frac{1}{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}P(a(X)=a^{\prime}|a^{\mathsf{c}}(X),C=S)}}$
@@
A.1.4 Unbiasedness of $\widehat{\tau}_{DR}$ Given One Correct Model
@@
$\mathbb{E}_{Y\sim P_{y}}$ $[\mathbb{E}_{X,X^{*}}[\widehat{\tau}_{DR}]]=\mathbb{E}_{Y\sim P_{y}}\Big{[}E_{X,X^{*}}\Big{[}\frac{1}{m}\sum_{j=1}^{m}\Big{[}\widehat{g}(1,a_{s}^{\mathsf{c}}(X^{*}_{j}))-\widehat{g}(0,a_{s}^{\mathsf{c}}(X^{*}_{j}))\Big{]}$
$\quad\quad\quad\quad\quad\quad\quad+\frac{1}{n}\sum_{i=1}^{n}\widehat{\gamma}(a(X_{i}),a_{s}^{\mathsf{c}}(X_{i}))(Y_{i}-\widehat{g}(a(X_{i}),a_{s}^{\mathsf{c}}(X_{i})))\Big{]}\Big{]}$
$=\frac{1}{m}\sum_{j=1}^{m}\mathbb{E}_{Y\sim P_{y}}[\mathbb{E}_{X^{*}}\left[\widehat{g}(1,a_{s}^{\mathsf{c}}(X^{*}_{j}))-\widehat{g}(0,a_{s}^{\mathsf{c}}(X^{*}_{j}))\right]]$
$\quad+\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}_{Y\sim P_{y}}[\mathbb{E}_{X}[\widehat{\gamma}(a(X_{i}),a_{s}^{\mathsf{c}}(X_{i}))(Y_{i}-\widehat{g}(a(X_{i}),a_{s}^{\mathsf{c}}(X_{i})))]]$
$=\frac{1}{m}\sum_{j=1}^{m}\mathbb{E}_{X^{*}}\left[\widehat{g}(1,a_{s}^{\mathsf{c}}(X^{*}_{j}))-\widehat{g}(0,a_{s}^{\mathsf{c}}(X^{*}_{j}))\right]$
$\quad+\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}_{Y\sim P_{y}}[\mathbb{E}_{X}[\mathbbm{1}\{a(X_{i})=1\}\widehat{\gamma}(1,a_{s}^{\mathsf{c}}(X_{i}))(Y_{i}-\widehat{g}(1,a_{s}^{\mathsf{c}}(X_{i})))$
$\quad+\mathbbm{1}\{a(X_{i})=0\}\widehat{\gamma}(0,a_{s}^{\mathsf{c}}(X_{i}))(Y_{i}-\widehat{g}(0,a_{s}^{\mathsf{c}}(X_{i})))]]$
@@
$\quad-\Bigg{(}\frac{1}{m}\sum_{j=1}^{m}\mathbb{E}_{X^{*}}[\widehat{g}(0,a_{s}^{\mathsf{c}}(X^{*}_{j}))]$
$\quad\quad\quad+\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}_{Y\sim P_{y}}\Bigg{[}\mathbb{E}_{X^{*}}\Bigg{[}\mathbbm{1}\{a(X_{i})=0\}\frac{P^{*}(a^{\mathsf{c}}(X_{i}))}{\widehat{P}(a_{s}^{\mathsf{c}}(X_{i}),a(X_{i})=0)}(Y_{i}-\widehat{g}(0,a_{s}^{\mathsf{c}}(X_{i}))\Bigg{]}\Bigg{]}\Bigg{)}$
@@
$=\mathbb{E}_{Y(\cdot)\sim\mathcal{G}}\Bigg{[}\mathbb{E}_{X}\Bigg{[}\sum_{x\in\mathcal{X}}\frac{P^{*}(a^{\mathsf{c}}(x))}{P(a^{\mathsf{c}}(x),a(x)=a^{\prime})}(Y(a^{\prime},a^{\mathsf{c}}(x))$
$\quad-\widehat{g}(a^{\prime},a_{s}^{\mathsf{c}}(x)))\mathbbm{1}\{a(x)=a^{\prime}\}\mathbbm{1}\{a^{\mathsf{c}}(X)=a^{\mathsf{c}}(x),a(X)=a(x)\}\Bigg{]}\Bigg{]}$
@@
$=\mathbb{E}_{Y(\cdot)\sim\mathcal{G}}\Bigg{[}\mathbb{E}_{X}\Bigg{[}\sum_{x\in\mathcal{X}}\mathbbm{1}\{a(x)=a^{\prime}\}\frac{P^{*}(a^{\mathsf{c}}(x))}{\widehat{P}(a_{s}^{\mathsf{c}}(x),a(x)=a^{\prime})}(Y(a^{\prime},a^{\mathsf{c}}(x))$
$\quad-g(a^{\prime},a^{\mathsf{c}}(x)))\mathbbm{1}\{a^{\mathsf{c}}(X)=a^{\mathsf{c}}(x),a(X)=a(x)\}\Bigg{]}\Bigg{]}$
$=\mathbb{E}_{Y(\cdot)\sim\mathcal{G}}\Bigg{[}\sum_{x\in\mathcal{X}}\frac{P^{*}(a^{\mathsf{c}}(x))}{\widehat{P}(a_{s}^{\mathsf{c}}(x),a(x)=a^{\prime})}(Y(a^{\prime},a^{\mathsf{c}}(x))$
$\quad-g(a^{\prime},a^{\mathsf{c}}(x)))\mathbb{E}_{X}\Bigg{[}\mathbbm{1}\{a(x)=a^{\prime}\}\mathbbm{1}\{a^{\mathsf{c}}(X)=a^{\mathsf{c}}(x),a(X)=a(x)\}\Bigg{]}\Bigg{]}$
@@
A.1.5 Confidence Intervals for $\widehat{\tau}_{DR}$
Consider data $(\widetilde{X}_{1},\dots,\widetilde{X}_{k})=(X_{1},\dots,X_{n},X^{*}_{1},\dots,X^{*}_{m})$, $(\widetilde{Y}_{1},\dots,\widetilde{Y}_{k})=(Y_{1},\dots,Y_{n},0,\dots,0)$. Then following standard procedures for doubly robust estimators (Kennedy, 2024), the estimator for the closed-form variance of $\widehat{\tau}_{DR}$ is derived from the influence function as follows.